OpenStudy (purple_pink):

A $7000 investment earns $1330 simple interest in two years. What is the annual interest rate? A. 1.90% B. 3.80% C. 9.50% D. 19.0%

OpenStudy (purple_pink):

@kittiwitti1

OpenStudy (kittiwitti1):

BOP! Hello ♪

OpenStudy (anonymous):

I literally just took a test on this

OpenStudy (kittiwitti1):

I believe this is one of those icky Principal formulas... xD

OpenStudy (anonymous):

one sec

OpenStudy (kittiwitti1):

*gives the mic to @CupcakeQueen10 * lol I haven't finished learning this subject yet I don't know for certain... ☻

OpenStudy (purple_pink):

okie np thanks

OpenStudy (anonymous):

I think its C let me explain

OpenStudy (anonymous):

I have this formula but since we already know the simple interest we need to do it in the inverse...

OpenStudy (anonymous):

so 1330/2=

OpenStudy (anonymous):

665 and 665/7000=

OpenStudy (anonymous):

0.095 and 0.095*100=

OpenStudy (anonymous):

9.50

OpenStudy (anonymous):

thats why i think its c

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

K i am back

OpenStudy (purple_pink):

mkay so A?

OpenStudy (anonymous):

no C

OpenStudy (anonymous):

sorry brb gotta check on my question

OpenStudy (purple_pink):

@amilliacortez123

OpenStudy (anonymous):

you don't get it or???

OpenStudy (purple_pink):

yes thank you! 2 more?

OpenStudy (anonymous):

maybe what the questions?

OpenStudy (purple_pink):

lemme make a new post

OpenStudy (anonymous):

ok :D